OpenStudy (anonymous):

If the point P(-3/5y) lies on the unit circle and (P) is in the second quadrant, what does (y) equal? HELP DX

OpenStudy (anonymous):

unit circle points satisfy this relation: \[\sqrt{x ^{2}+y ^{2}} =1\] so: \[\sqrt{(3/5)^{2}+y ^{2}} = 1\] solve this equation for y. You will get 2 values of y which are bouh correct

OpenStudy (anonymous):

@jCortez93

OpenStudy (anonymous):

if you have problem solving it, tell me

OpenStudy (anonymous):

How do you solve it?

OpenStudy (anonymous):

square bouth sides and you get: \[3/5^{2}+y ^{2} =1\] so: \[y= \pm \sqrt{1-9/25}=\pm \sqrt{16/25} =\pm 4/5\]

OpenStudy (anonymous):

got it?